# Efficient Approximate Degenerate Ordered Statistics Decoding for Quantum Codes via Reliable Subset Reduction

### v3

@@ Abstract
Efficient and scalable decoding of quantum codes is essential for high-performance quantum error correction. In this work, we introduce Reliable Subset Reduction (RSR), a reliability-driven preprocessing framework that leverages belief propagation (BP) statistics to identify and remove highly reliable qubits, substantially reducing the effective problem size. Additionally, we identify a degeneracy condition that allows high-order OSD to be simplified to order-0 OSD. By integrating these techniques, we present an ADOSD algorithm that significantly improves OSD efficiency. Our BP+RSR+ADOSD framework extends naturally to circuit-level noise and can handle large-scale codes with more than $10^{4}$ error variables. Through extensive simulations, we demonstrate improved performance over MWPM and Localized Statistics Decoding for a variety of CSS and non-CSS codes under the code-capacity noise model, and for rotated surface codes under realistic circuit-level noise. At low physical error rates, RSR reduces the effective problem size to less than 5%, enabling higher-order OSD with accelerated runtime. These results highlight the practical efficiency and broad applicability of the BP+ADOSD framework for both theoretical and realistic quantum error correction scenarios.
@@ I Introduction
In this paper, we investigate postprocessing techniques for enhancing the performance of BP. When BP fails to converge to a valid global solution, it still provides rich statistical information about the error configuration. In particular, motivated by density evolution arguments for sparse LDPC codes [58, 59], most error variables are typically resolved with high confidence, except for a small number of unsettled variables associated with trapping sets [30]. Consequently, a large fraction of variables can be regarded as already determined and effectively part of the solution.
@@
Beyond improved decoding performance, a central outcome of our simulations is the strong reduction effect produced by reliable subset reduction (RSR): at low physical error rates, RSR typically reduces the effective decoding problem size by one to two orders of magnitude (e.g., below $5\%$ of the original size for $\epsilon\leq 0.005$, and below $1\%$ for $\epsilon=0.001$ in surface-code DEM problems). This dramatic dimensionality reduction enables higher-order OSD to be applied efficiently even for problems with more than $10^{4}$ error variables, a regime that is far beyond the reach of conventional OSD-based postprocessing. As a result, our MBP–RSR–ADOSD approach achieves strong threshold performance for topological codes and substantially improved logical error rates in the low-error regime.
@@ III The Reliable Subset Reduction Method
In this section, we estimate the expected size of these unsettled error variables using density evolution [58, 59].
III-A Density-Evolution-Guided Reliability Heuristic
We consider a family of regular LDPC codes over a quaternary symmetric channel and use the syndrome-based Gallager-A algorithm [15] as a heuristic approximation of quantum stabilizer decoding under depolarizing noise. Density evolution is then employed to characterize the behavior of this iterative decoder.
Gallager-A is a hard-decision message-passing algorithm defined on a Tanner graph. In our setting, variable nodes $v_{j}$ represent Pauli errors on physical qubits, while check nodes $c_{i}$ represent stabilizer generators. Since a Pauli operator either commutes or anticommutes with a stabilizer, all messages are binary: $1$ indicates anti-commutation and 0 indicates commutation.
We consider a syndrome-based quaternary–binary variant of Gallager-A suitable for CSS-type stabilizer codes, where each stabilizer contains only $X$ or $Z$ operators. Given a check matrix $\mathbf{H}$} and syndrome ${\mathbf{s}}=(s_{1},\dots,s_{m})$, the decoding proceeds as follows.
  1. 1.
    Initialization: Set estimated error $\hat{E}_{j}=I$ for all qubits and initialize all variable-to-check messages
    $m_{v_{j}\to c_{i}}^{(0)}=0.$
  2. 2.
    Check update: At iteration $\ell$, each check node computes a recommendation message
    $m_{c_{i}\to v_{j}}^{(\ell)}=s_{i}\oplus\bigoplus_{k\in{\cal I}(c_{i})\setminus\{j\}}m_{v_{k}\to c_{i}}^{(\ell)}.$
  3. 3.
    Variable update: Each variable node $v_{j}$ updates its estimate according to the unanimous recommendations of its neighboring checks:
    • $X$-flip: if all incoming messages from $Z$-type checks are $1$, set $\hat{E}_{j}\leftarrow\hat{E}_{j}X$;
    • $Z$-flip: if all incoming messages from $X$-type checks are $1$, set $\hat{E}_{j}\leftarrow\hat{E}_{j}Z$.
    (If both updates occur, a $Y$ correction is effectively applied up to a global phase.)
  4. 4.
    Message update: Set $m_{v_{j}\to c_{i}}^{(\ell+1)}=1$ if $\hat{E}_{j}$ anticommutes with the stabilizer of $c_{i}$, and 0 otherwise.
  5. 5.
    Termination: The algorithm stops if the current estimate $\hat{\mathbf{E}}$ is valid.
For a CSS-type stabilizer code, $X$-type and $Z$-type errors are processed independently by the modified Gallager-A algorithm. The correlations between $X$ and $Z$ errors induced by the depolarizing channel are therefore neglected at this stage. Accordingly, we focus on the analysis of $X$-error decoding, which reduces to a classical binary decoding problem: estimating an $n$-bit vector ${\mathbf{e}}=(e_{1},\dots,e_{n})\in\{0,1\}^{n}$ from its syndrome.
Suppose that the $Z$-type stabilizers define a $(d_{c},d_{v})$-regular LDPC code, where each check node is connected to $d_{c}$ variable nodes and each variable node is connected to $d_{v}$ check nodes.
Proposition 2.
[Heuristic Density Evolution Analysis] Consider a quantum code of length $n$ under depolarizing noise with rate $\epsilon$. Assume that the Tanner graph induced by the $Z$ stabilizers is locally tree-like and that decoding is performed using the modified Gallager-A algorithm. Let $x^{(\ell)}$ denote the probability that a variable-to-check message is incorrect at iteration $\ell$, and let $y^{(\ell)}$ denote the probability that a check-to-variable message is incorrect. Then the message error probabilities evolve according to
$y^{(\ell)}$ $=\frac{1-(1-2x^{(\ell)})^{d_{c}-1}}{2},$ (6)
$x^{(\ell+1)}$ $=p_{a}\big[1-(1-y^{(\ell)})^{d_{v}-1}\big]+(1-p_{a})(y^{(\ell)})^{d_{v}-1},$ (7)
where $p_{a}=\frac{2\epsilon}{3}$ is the probability that a Pauli error anticommutes with a $Z$-type stabilizer, and $x^{(0)}=p_{a}$.
The recursion (6)–(7) follows from the standard density-evolution analysis of Gallager-A decoding for regular LDPC codes over the binary symmetric channel [15, 58, 59].
At error rates below the heuristic threshold, density evolution predicts that most variable nodes converge rapidly, leaving a fraction proportional to the channel error probability $\epsilon$ unresolved. The expected number of erroneous variable nodes at iteration $\ell$ is $n\cdot x^{(\ell)}$, and after an average convergence time $\ell_{\mathrm{succ}}$ the residual error set has size roughly $n\cdot x^{(\ell_{\mathrm{succ}})}$. For finite-length codes, cycles in the Tanner graph cause deviations from the infinite-length density evolution prediction. As $\epsilon$ decreases, $x^{(\ell_{\mathrm{succ}})}$ rapidly vanishes, leaving extremely sparse residual errors.
Unlike classical LDPC ensembles, stabilizer Tanner graphs inherently contain many short cycles due to commutation constraints. As a result, the above density-evolution analysis should be interpreted as a heuristic guideline rather than a rigorous asymptotic prediction. Although one could in principle extend DE to full belief propagation with real-valued LLRs, the Gallager-A approximation already captures the main qualitative behavior of the decoding process. Nonetheless, our heuristic results serve as a useful reference for BP algorithms that exploit memory effects to mitigate the impact of short cycles [32, 77].
III-B Reliable subset reduction for the syndrome decoding problem
@@
Motivated by density evolution, we expect that after a few BP iterations, most variables have already converged to highly reliable values, while only a small subset remains ambiguous. We denote by $R\subset\{e_{1},\dots,e_{2n}\}$ a subset of reliable variables, whose values are assumed to be correct and fixed. The remaining variables are treated as unknowns.
@@
III-C Reliability order based on BP4
@@ IV Order-statistic decoding based on quaternary belief propagation
Once BP fails to produce a valid error estimate, combinatorial post-processing techniques are required. These techniques operate on the residual errors left by BP and can include methods such as OSD, LSD, or other syndrome-based decoders. In this section, we first review the basic OSD procedure, which relies on the reliability order defined in Section III-C. When combined with the RSR method introduced in Section III-B, OSD can operate on a reduced system of variables, thereby lowering computational complexity.
In the context of OSD, a critical step involves identifying and sorting the accurate and reliable coordinates to establish $n+k$ linearly independent bits. Then, we can use these coordinates to generate a valid error that matches the syndrome. To accomplish this, we use the reliability orders defined in Section III-C.
@@
4:Let $\mu(\pi(\tilde{{\mathbf{e}}}))=\left[\tilde{{\mathbf{e}}}^{U}~~\tilde{{\mathbf{e}}}^{R}\right]$, where $\tilde{{\mathbf{e}}}^{U}$ represents the independent and unreliable subset of nkn-k bits and $\tilde{{\mathbf{e}}}^{R}$ is the reliable subset of $n+k$ bits. Use the reliable subset $\tilde{{\mathbf{e}}}^{R}$ and the modified parity-check condition 12 to select feasible unreliable bits
@@ IV-C Reliable Subset Reduction combined with OSD
As introduced in Section III-B, the RSR procedure removes highly reliable error variables from the parity-check constraints, producing a reduced linear system. This reduced system can then be solved using OSD or any other suitable combinatorial decoder to estimate the remaining unreliable variables.
@@ VII-B1 Bivariate bicycle codes
TABLE III: Comparison of the time consumption for surface codes at LER about $10^{-6}$ with $\alpha=1.0$, $T=100$, and $\theta=0.999995$.
code $\epsilon$ MBP4 time per MBP4 time per time per average iterations ADOSD4 time
[…]
11 1.7% 21.46% 0.01464 0.0272 0.667 1.002 2
13 2.5% 35.97% 0.0212 0.0469 1.342 1.79 2.21
15 3.3% 51.47% 0.0287 0.0506 2.155 2.41 1.76
$^{\dagger}:$ the time per ADOSD4 operation divided by the time per MBP4 iteration.
VII-B3 Time Consumption of ADOSD4 vs. OSD4-2 and BP
[…]
It can be observed that the time consumption of ADOSD4 is approximately 2% to 4% of that of OSD4-2, with similar results observed for other quantum codes in this study. In summary, ADOSD4 is at least 25 times faster than OSD4-2 at an LER of around $10^{-6}$.
Table III also lists the time consumption of ADOSD4 in terms of MBP4 iterations. Specifically, the average time consumption per codeword for ADOSD4 is approximately equivalent to 1.76 to 2.21 MBP4 iterations. This shows that the time consumption for ADOSD4 is roughly equivalent to 1 to 2 successful MBP4 executions.
@@ VII-C Circuit-Level Surface Code Decoding
Note that decoders which treat $X$ and $Z$ errors separately, such as the standard MWPM decoder, do not make use of both types of detector variables. For comparison, we therefore also generate decoding problems in which the $X$-type detector variables are removed. The parameters of the resulting DEM decoding problems are summarized in Table IV.
@@
TABLE IV: Parameters of the DEM decoding problems for rotated surface codes. Subtable (a) shows the default setting with both $X$ and $Z$ detector variables, and subtable (b) shows the modified setting with only $Z$ detector variables.
@@ VII-C1 Decoding performance
For reference, as shown in Ref. [90], RelayBP performs worse than MWPM for rotated surface codes. We also remark that MWPM can be adapted to exploit $X/Z$ correlations [91, 92, 93].
@@
VII-C2 Decoding statistics
We summarize the decoding statistics for circuit-level surface-code decoding of distance $9$ without $X$-type detector variables in Table V. The row “Samples” reports the total number of Monte Carlo shots, while “OSD calls” counts the number of samples for which BP does not converge to a valid solution and OSD is therefore invoked.
The row “Average length after RSR” shows that the effective code length after reduction is below $5\%$ of the original for $\epsilon\leq 0.005$, and below $1\%$ for $\epsilon=0.001$, demonstrating the strong reduction capability of RSR.
From the statistics, we observe no stage-1 failures and only a small number of stage-2 failures. Even if all stage-2 failures are conservatively counted as logical errors, the overall logical error rate is only marginally affected.
@@ VII-C2 Decoding statistics
“Time per BP iteration” denotes the average time of a single BP iteration, averaged over all runs. “Time per OSD” is the average time of one OSD call, including both ADOSD and the fallback regular OSD when RSR fails. We observe that the time per BP iteration is essentially independent of the depolarizing rate $\epsilon$. In contrast, the OSD time decreases significantly as $\epsilon$ becomes smaller. At $\epsilon=0.01$, one OSD call is equivalent to roughly 250 BP iterations, while at $\epsilon=0.001$ it drops to about 18 BP iterations.
@@
Finally, we remark that the current OSD runtimes do not yet fully reflect the potential speedup enabled by RSR. Our implementation embeds the binary DEM decoding problem into a quaternary formalism, which is not optimized for this task and can be further accelerated by a dedicated binary implementation.
TABLE V: Statistics of circuit-level surface decoding of distance $9$ without $X$ detector variables. $N=1945$.
@@
Samples 100,000 100,000 115990 1311328 9943799 10,000,000
OSD calls 99966 96895 91954 483438 1632156 389602
Average length after RSR 229.76 113.37 62.64 32.69 22.67 14.99
relative code length 0.12 0.06 0.03 0.02 0.012 0.008
@@
Stage-2 Failures 15 43 109 1142 3927 916
OSD-0 0 0 0 2 107 219
Avg iterations w/o OSD 5.824 5.691 5.144 4.525 4.059 3.353
Time per BP iteration (ms) 0.1565 0.1498 0.1289 0.1248 0.1219 0.1185
Time per OSD (ms) 37.65 27.04 20.71 19.90 8.632 1.987

### v4

@@ Abstract
Efficient and scalable decoding of quantum codes is essential for high-performance quantum error correction. In this work, we introduce Reliable Subset Reduction (RSR), a reliability-driven preprocessing framework that leverages belief propagation (BP) statistics to identify and remove highly reliable qubits, substantially reducing the effective problem size. Additionally, we identify a degeneracy condition that allows high-order OSD to be simplified to order-0 OSD. By integrating these techniques, we present an ADOSD algorithm that significantly improves OSD efficiency. Our BP+RSR+ADOSD framework extends naturally to circuit-level noise and can handle large-scale codes with more than $10^{4}$ error variables. Through extensive simulations, we demonstrate improved performance over MWPM and Localized Statistics Decoding for a variety of CSS and non-CSS codes under the code-capacity noise model, and for rotated surface codes under realistic circuit-level noise. At low physical error rates, RSR reduces the effective problem size to as little as 1% (e.g., for $\epsilon=0.001$ in surface-code DEM), enabling higher-order OSD with drastically reduced computational complexity. These results highlight the practical efficiency and broad applicability of the BP+ADOSD framework for both theoretical and realistic quantum error correction scenarios.
@@ I Introduction
In this paper, we investigate postprocessing techniques for enhancing the performance of BP. When BP fails to converge to a valid global solution, it still provides rich statistical information about the error configuration. In particular, observing the convergence behavior of belief propagation on sparse graphs, most error variables are typically resolved with high confidence, except for a small number of unsettled variables associated with trapping sets [30]. Consequently, a large fraction of variables can be regarded as already determined and effectively part of the solution.
@@
Beyond improved decoding performance, a central outcome of our simulations is the strong reduction effect produced by reliable subset reduction (RSR): at low physical error rates, RSR typically reduces the effective decoding problem size by one to two orders of magnitude (e.g., below $4\%$ of the original size for $\epsilon\leq 0.005$, and below $1\%$ for $\epsilon=0.001$ in surface-code DEM problems). This dramatic dimensionality reduction enables higher-order OSD to be applied efficiently even for problems with more than $10^{4}$ error variables, a regime that is far beyond the reach of conventional OSD-based postprocessing. As a result, our MBP–RSR–ADOSD approach achieves strong threshold performance for topological codes and substantially improved logical error rates in the low-error regime. Crucially, these results are achieved under a strict budget of at most 10 BP iterations. In practice, when MBP4 converges, it typically does so within 3 to 6 iterations.
@@ III The Reliable Subset Reduction Method
III-A Reliable subset reduction for the syndrome decoding problem
@@
We expect that after a few BP iterations, most variables have already converged to highly reliable values, while only a small subset remains ambiguous. We denote by $R\subset\{e_{1},\dots,e_{2n}\}$ a subset of reliable variables, whose values are assumed to be correct and fixed. The remaining variables are treated as unknowns.
@@
III-B Reliability order based on BP4
@@ IV Order-statistic decoding based on quaternary belief propagation
Once BP fails to produce a valid error estimate, combinatorial post-processing techniques are required. These techniques operate on the residual errors left by BP and can include methods such as OSD, LSD, or other syndrome-based decoders. In this section, we first review the basic OSD procedure, which relies on the reliability order defined in Section III-B. When combined with the RSR method introduced in Section III-A, OSD can operate on a reduced system of variables, thereby lowering computational complexity.
In the context of OSD, a critical step involves identifying and sorting the accurate and reliable coordinates to establish $n+k$ linearly independent bits. Then, we can use these coordinates to generate a valid error that matches the syndrome. To accomplish this, we use the reliability orders defined in Section III-B.
@@
4:Let $\mu(\pi(\tilde{{\mathbf{e}}}))=\left[\tilde{{\mathbf{e}}}^{U}~~\tilde{{\mathbf{e}}}^{R}\right]$, where $\tilde{{\mathbf{e}}}^{U}$ represents the independent and unreliable subset of nkn-k bits and $\tilde{{\mathbf{e}}}^{R}$ is the reliable subset of $n+k$ bits. Use the reliable subset $\tilde{{\mathbf{e}}}^{R}$ and the modified parity-check condition 10 to select feasible unreliable bits
@@ IV-C Reliable Subset Reduction combined with OSD
As introduced in Section III-A, the RSR procedure removes highly reliable error variables from the parity-check constraints, producing a reduced linear system. This reduced system can then be solved using OSD or any other suitable combinatorial decoder to estimate the remaining unreliable variables.
@@ VII-C Circuit-Level Surface Code Decoding
Note that decoders which treat $X$ and $Z$ errors separately, such as the standard MWPM decoder, do not make use of both types of detector variables. For comparison, we therefore also generate decoding problems in which the $X$-type detector variables are removed. The parameters of the resulting DEM decoding problems are summarized in Table III.
@@
TABLE III: Parameters of the DEM decoding problems for rotated surface codes. Subtable (a) shows the default setting with both $X$ and $Z$ detector variables, and subtable (b) shows the modified setting with only $Z$ detector variables.
@@
For reference, as shown in [88], RelayBP performs worse than MWPM for rotated surface codes. We also remark that MWPM can be adapted to exploit $X/Z$ correlations [89, 90, 91].
@@
VII-D Decoding statistics
We summarize the decoding statistics for circuit-level surface-code decoding of distance $9$ without $X$-type detector variables in Table IV, using $\alpha=1.5$, $T=10$, and $\theta=0.99$. The row “Shots” reports the total number of Monte Carlo shots, while “OSD calls/Shots” is the percentage of shots for which BP does not converge to a valid solution and OSD is therefore invoked. The row “Average length after RSR” shows that the effective code length after reduction is below $4\%$ of the original for $\epsilon\leq 0.005$, and below $1\%$ for $\epsilon=0.001$, demonstrating the strong reduction capability of RSR.
From the statistics, we observe no stage-1 failures and only a small percentage of stage-2 failures. Even if all stage-2 failures are conservatively counted as logical errors, the overall logical error rate is only marginally affected.
@@ VII-D Decoding statistics
Finally, we remark that the current software runtimes do not yet fully reflect the intrinsic speedup enabled by RSR. Our current research-oriented implementation embeds the binary DEM decoding problem into a quaternary formalism, which incurs unnecessary computational overhead. A dedicated binary implementation is expected to further reduce the wall-clock time.
TABLE IV: Statistics of circuit-level surface decoding ($d=9,N=1945$) without $X$ detector variables. Parameters: $\alpha=1.5,T=10,\theta=0.99$. $n_{red}$ denotes the reduced variable length by RSR.
@@
Shots $10^{6}$ $10^{6}$ $10^{6}$ $10^{6}$ $10^{7}$ $10^{7}$
OSD calls / Shots $99.97\%$ $96.85\%$ $79.12\%$ $36.88\%$ $16.41\%$ $3.90\%$
Average length after RSR ($n_{red}$) 229.53 113.27 62.34 32.85 22.67 14.99
Relative code length ($n_{red}/N$) 0.118 0.058 0.032 0.017 0.012 0.008
@@
Stage-2 Failures / Shots $1.19\times 10^{-4}$ $4.33\times 10^{-4}$ $61\times 10^{-4} 8.8$$88\times 10^{-4}$ $3.95\times 10^{-4}$ $9.16\times 10^{-5}$
Avg. BP iterations (converged) 6.21 5.67 5.15 4.53 4.07 3.50
The order-of-magnitude reduction in the residual problem size ensures that the computational cost of the ADOSD stage remains comparable to only a few BP iterations. This low-complexity feature is consistent across both code-capacity and circuit-level noise models, facilitating the implementation of high-order OSD in large-scale quantum error correction systems.